# iclr2025_conference.bib

### v3

@@ Appendix B Explanation of the $N_{G}$’s importance in fair comparison
The importance of $N_{G}$ lies in its impact on the non-linear expectation of uniquely memorized images. Comparing UMS across different $N_{G}$ values is flawed because varying $N_{G}$ leads to different outcomes, underscoring the need for consistent $N_{G}$ values to ensure fair comparisons. In carlini2023extracting, the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images. Our theoretical analysis accurately aligns with the behavior observed in the experimental data in Fig 5.
@@ Appendix E Pseudocode for SIDE method
The following Algorithm 1 outlines the detailed steps of the SIDE method for extracting training data from unconditional diffusion models. This algorithm combines the construction of implicit informative labels, the training of a time-dependent classifier, and the conditional generation process to effectively extract valuable training samples.
@@
20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using equation (151)
@@ Appendix F SSCD-related metrics
In our evaluation of the proposed method, it is essential to include a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS (Average Matching Similarity) and UMS (Unconditional Matching Similarity). While 95th percentile SSCD metric has its limitations, it still serves as a useful reference point for assessing the relative similarity between extracted images and the original training dataset.
@@
In addition, we report detailed metrics in the context of our experiments on the CelebA-HQ-FI dataset, which further illustrates the effectiveness of our approach. These results underscore the significance of incorporating SSCD-related scores to complement AMS and UMS in providing a more nuanced understanding of the similarities between generated and training images.
@@
Table 4: Generate Training Epoch: 3000 Dataset: CelebA-HQ-FI Generate Nums Per $\lambda$: 50000. The AMS and UMS is measured on Mid Similarity
AMS(%) UMS(%) Top 0.1% Top 0.5% Top 1.0% Top 5.0% Top 10.0%
$\lambda$
0 0.596 0.328 0.604 0.544 0.518 0.463 0.440
1 0.588 0.312 0.596 0.540 0.517 0.463 0.440
2 0.640 0.350 0.591 0.541 0.518 0.465 0.441
3 0.764 0.386 0.594 0.549 0.525 0.470 0.446
4 0.850 0.390 0.604 0.553 0.529 0.473 0.448
5 0.952 0.436 0.596 0.551 0.530 0.476 0.451
6 1.092 0.414 0.611 0.560 0.536 0.480 0.454
7 1.110 0.446 0.607 0.562 0.539 0.482 0.457
8 1.148 0.444 0.618 0.566 0.542 0.484 0.458
9 1.274 0.478 0.615 0.567 0.544 0.485 0.459
10 1.338 0.444 0.613 0.569 0.546 0.487 0.461
11 1.292 0.454 0.604 0.562 0.541 0.486 0.460
12 1.262 0.406 0.617 0.567 0.544 0.486 0.460
13 1.390 0.432 0.617 0.569 0.546 0.489 0.462
14 1.232 0.384 0.613 0.567 0.544 0.485 0.459
15 1.516 0.462 0.616 0.570 0.548 0.490 0.463
16 1.280 0.390 0.612 0.566 0.543 0.487 0.461
17 1.282 0.386 0.605 0.561 0.541 0.486 0.460
18 1.330 0.374 0.616 0.569 0.545 0.488 0.461
19 1.204 0.354 0.612 0.564 0.541 0.485 0.460
20 1.178 0.358 0.603 0.559 0.538 0.483 0.458
21 1.172 0.342 0.617 0.566 0.542 0.484 0.459
22 1.208 0.368 0.602 0.560 0.539 0.485 0.459
23 1.286 0.302 0.607 0.561 0.540 0.485 0.459
24 1.244 0.352 0.597 0.558 0.538 0.484 0.458
25 1.198 0.340 0.599 0.560 0.538 0.483 0.458
26 1.220 0.338 0.601 0.559 0.539 0.483 0.458
27 1.128 0.320 0.608 0.561 0.538 0.483 0.457
28 1.102 0.314 0.604 0.556 0.534 0.481 0.456
29 1.034 0.290 0.595 0.556 0.534 0.481 0.456
30 1.026 0.326 0.602 0.557 0.535 0.480 0.455
31 1.020 0.268 0.591 0.551 0.531 0.479 0.455
32 1.054 0.282 0.593 0.551 0.531 0.479 0.455
33 1.106 0.306 0.600 0.555 0.535 0.481 0.456
34 1.062 0.288 0.582 0.547 0.529 0.479 0.454
35 0.922 0.266 0.587 0.547 0.527 0.477 0.453
36 0.874 0.260 0.585 0.545 0.525 0.477 0.453
37 0.964 0.258 0.589 0.549 0.528 0.477 0.452
38 0.888 0.246 0.582 0.543 0.524 0.475 0.452
39 0.940 0.274 0.587 0.548 0.528 0.476 0.452
40 0.808 0.234 0.587 0.544 0.524 0.474 0.451
41 0.870 0.252 0.582 0.543 0.524 0.476 0.452
42 0.872 0.238 0.584 0.543 0.523 0.475 0.451
43 0.856 0.244 0.584 0.545 0.525 0.475 0.451
44 0.796 0.212 0.578 0.540 0.521 0.473 0.449
45 0.770 0.242 0.580 0.538 0.519 0.472 0.449
46 0.774 0.218 0.580 0.540 0.521 0.472 0.448
47 0.754 0.214 0.581 0.542 0.521 0.471 0.448
48 0.716 0.218 0.572 0.536 0.518 0.471 0.448
49 0.694 0.216 0.570 0.533 0.515 0.469 0.446
50 0.728 0.204 0.576 0.535 0.518 0.471 0.447

### v5

@@ Appendix B Explanation UMS’s importance in fair comparison
The importance of $N_{G}$ lies in its impact on the non-linear expectation of uniquely memorized images. Comparing UMS across different $N_{G}$ values is flawed because varying $N_{G}$ leads to different outcomes, underscoring the need for consistent $N_{G}$ values to ensure fair comparisons. In [5], the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images. Our theoretical analysis accurately aligns with behavior observed in the experimental data in Fig 6.
Appendix C Refinement ResNet block (Fig. 7)
@@ Appendix E Pseudocode for SIDE method
The following Algorithm 1 outlines the detailed steps of the SIDE method for extracting training data from unconditional diffusion models. This algorithm combines the construction of implicit informative labels, the training of a time-dependent classifier, and the conditional generation process to extract valuable training samples effectively.
@@
20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using Eq. (22)
@@ Appendix F SSCD-related metrics
In our evaluation of the proposed method, it is essential to include a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS and UMS. While 95th percentile SSCD metric has its limitations, it still serves as a useful reference point for assessing the relative similarity between extracted images and the original training dataset.
@@
In addition, we report detailed metrics in the context of our experiments on the CelebA-HQ-FI dataset, further illustrating our approach’s effectiveness.
@@
Appendix G Extended SIDE for Text-to-Image Diffusion Models
We extend the SIDE method for a text-to-image diffusion model in a fine-tuning scenario. Instead of using a classifier as a surrogate condition to extract training data, we use random strings as surrogate conditions. The attack setting is simple and practical: the attacker can access the fine-tuned model and modify the text inputs in the text-to-image pairs. Notably, the attacker does not have access to the training data. We name this method as Extended SIDE.
G.1 Attack Setting
The attack setting for the Extended SIDE method involves two primary components: the victim’s fine-tuning process and the attacker’s access to the fine-tuned model. Below, we provide a detailed description of each component.
  • Victim Fine-Tuning:
    The victim fine-tunes a text-to-image diffusion model, denoted as $p_{\theta}(\bm{x}\mid s)$, using a dataset of image-text pairs $\mathcal{D}_{train}=\{(s_{i},\bm{x}_{i})\}_{i=1}^{N}$. In this context:
    • $s_{i}\in\mathcal{S}_{s}$ are random strings crafted by the attacker to serve as surrogate text prompts.
    • $\bm{x}_{i}\in\mathbb{R}^{d}$ represents the corresponding images associated with each random string.
    The fine-tuning objective is to optimize the model parameters $\theta$ to minimize the loss function over the training data:
    $\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(s,\bm{x})\sim\mathcal{D}_{train}}\left[\mathcal{L}(p_{\theta}(\bm{x}\mid s),\bm{x})\right]$ (28)
    where $\mathcal{L}$ denotes the loss function, such as mean squared error (MSE) or cross-entropy loss. The use of random strings as surrogate conditions aims to induce memorization within the model, causing it to overfit specific image-text associations during the fine-tuning process.
  • Attacker’s Access:
    After the fine-tuning phase, the attacker gains access to the fine-tuned model $p_{\theta}(\bm{x}\mid s)$. This access allows the attacker to perform the following actions:
    • Input Manipulation: The attacker can input any text prompt $s\in\mathcal{S}_{s}$ into the model to generate corresponding images:
      $\bm{x}\sim p_{\theta}(\bm{x}\mid s)$ (29)
    • Data Extraction: By systematically querying the model with the same set of random strings $\mathcal{S}_{s}$, the attacker can generate multiple images $\{\bm{x}_{i,j}\}_{j=1}^{N_{G}}$ for each string $s_{i}$. The goal is to identify images that closely resemble the memorized training samples:
      $\bm{x}_{i,j}\sim p_{\theta}(\bm{x}\mid s_{i}),\quad j=1,\dots,N_{G}$ (30)
    • Memorization Detection: The attacker analyzes the consistency and similarity among the generated images for each random string $s_{i}$. A low variance $\sigma_{s_{i}}^{2}$ among the samples indicates potential memorization:
      $\sigma_{s_{i}}^{2}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i,j}-\overline{\bm{x}}_{s_{i}}\|^{2}$ (31)
      where $\overline{\bm{x}}_{s_{i}}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i,j}$ is the mean image for string $s_{i}$. If $\sigma_{s_{i}}^{2}$ falls below a predefined threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s_{i}}$ as a potential memorized training sample.
  • Assumptions:
    • Knowledge of Surrogate Strings: The attacker is aware of the set of random strings $\mathcal{S}_{s}$ used during the victim’s fine-tuning process.
    • Model Query Capability: The attacker can perform a large number of queries to the fine-tuned model with each random string $s_{i}$ to generate sufficient samples for variance analysis.
    • No Access to Training Data: The attacker does not have direct access to the victim’s training dataset $\mathcal{D}_{train}$.
  • Attack Goals:
    • Data Reconstruction: To reconstruct or approximate the training images $\bm{x}_{i}$ by exploiting the model’s memorization of the random string-image associations.
    • Privacy Breach: To demonstrate the potential privacy risks associated with fine-tuned diffusion models by extracting sensitive or proprietary training data.
G.2 Extended SIDE Method
The Extended SIDE method leverages random strings as surrogate conditions to directly extract training images from a fine-tuned diffusion model. Unlike the original SIDE method, which relies on a classifier to guide the extraction process, this approach simplifies the attack by using random strings to induce and exploit memorization within the model.
Algorithm 2 Extended SIDE Method for Extracting Training Data Using Random Strings
1:Input:
  • Fine-tuned diffusion model $p_{\theta}(\bm{x}\mid s)$
  • Set of random strings $\mathcal{S}_{s}$
  • Number of samples per string $N_{G}$
  • Variance threshold $\tau$
2:Output: Extracted data $\mathcal{D}_{\text{extracted}}$
3:Initialize $\mathcal{D}_{\text{extracted}}\leftarrow\emptyset$
4:for each $s\in\mathcal{S}_{s}$ do
5:     Generate samples $\{\bm{x}_{i}^{(j)}\}_{j=1}^{N_{G}}$ where $\bm{x}_{i}^{(j)}\sim p_{\theta}(\bm{x}\mid s)$
6:     Compute mean image:
$\overline{\bm{x}}_{s}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i}^{(j)}$ (32)
7:     Compute sample variance:
$\sigma_{s}^{2}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i}^{(j)}-\overline{\bm{x}}_{s}\|^{2}$ (33)
8:     if $\sigma_{s}^{2}<\tau$ then
9:         $\mathcal{D}_{\text{extracted}}\leftarrow\mathcal{D}_{\text{extracted}}\cup\{\overline{\bm{x}}_{s}\}$
10:     end if
11:end for
12:return $\mathcal{D}_{\text{extracted}}$
The Extended SIDE method operates on the principle that random strings, when used as surrogate conditions during fine-tuning, become tightly coupled with specific training images. This follows the results of our theoretical understanding in Theorem 1. The attacker exploits this behavior to extract training data by performing the following steps:
  1. 1.
    Selection of Random Strings: The attacker identifies or generates a set of random strings $\mathcal{S}_{s}$ that were used during the victim’s fine-tuning process. These strings act as unique keys that the model associates with specific images.
  2. 2.
    Image Generation: For each random string $s\in\mathcal{S}_{s}$, the attacker generates $N_{G}$ images by querying the fine-tuned model:
    $\bm{x}_{i}^{(j)}\sim p_{\theta}(\bm{x}\mid s),\quad j=1,\dots,N_{G}$ (34)
    The repetition of image generation for the same string allows the attacker to assess the consistency of the outputs.
  3. 3.
    Consistency Analysis: The attacker computes the mean image $\overline{\bm{x}}_{s}$ and the sample variance $\sigma_{s}^{2}$ for the generated images associated with each string $s$:
    $\overline{\bm{x}}_{s}$ $=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i}^{(j)}$ (35)
    $\sigma_{s}^{2}$ $=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i}^{(j)}-\overline{\bm{x}}_{s}\|^{2}$ (36)
    A low variance $\sigma_{s}^{2}<\tau$ indicates that the generated images are highly similar, suggesting that the model has memorized the association between the string $s$ and a specific training image $\bm{x}_{s}$.
  4. 4.
    Data Extraction: If the variance $\sigma_{s}^{2}$ is below the threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s}$ as an extracted training image and adds it to the dataset $\mathcal{D}_{\text{extracted}}$:
    $\mathcal{D}_{\text{extracted}}\leftarrow\mathcal{D}_{\text{extracted}}\cup\{\overline{\bm{x}}_{s}\}$ (37)
This method effectively leverages the deterministic mapping between random strings and training images induced by memorization, allowing the attacker to reconstruct the training dataset without direct access.
G.3 Algorithmic Implementation
The Extended SIDE method is formalized in Algorithm 2. The algorithm outlines the steps the attacker takes to generate images, compute the variance, and extract the training data based on the consistency of generated images. Since the attacker cannot compute similarity to the training data, they detect memorization by analyzing the diversity of generated images conditioned on the same random string $s$. The attacker identifies memorization by detecting low variance in the images generated from the same random string $s$. A low variance $\sigma_{s}^{2}$ indicates that the model consistently generates similar images for $s$, suggesting that it has memorized specific training data associated with $s$. The Extended SIDE is shown in Algorithm 2
G.4 Experiments
To verify the effectiveness of the Extended SIDE, we conduct experiments on the subset of the Laion-2B dataset, finetuning the Stable Diffusion v1.5 using the LORA[18] method. We use the following metrics to evaluate the attack results: mean SSCD (M-SSCD), AMS (mid-similarity) and LIPIPS. The results are shown in Tab. 4. We use the generated images of original prompts as the Baseline method. The extracted examples are shown in Figure 8. As the results show, our method could effectively extract training data from fine-tuned DPMs.
Table 4: Performance Comparison for Extended SIDE method
Dataset Method M-SSCD AMS LIPIPS
Subset of Laion-2B Baseline 0.028 0.000 0.892
Extended SIDE(ours) 0.467 0.672 0.809
As presented in Tab. 4, the Extended SIDE method outperforms the baseline across all evaluation metrics:
  • M-SSCD: The SIDE method achieves an M-SSCD of 0.467 compared to the baseline’s 0.028. This significant increase indicates that SIDE extracts images with higher structural similarity to the training data, demonstrating effective memorization.
  • AMS: SIDE records an AMS of 0.672, markedly higher than the baseline’s 0.000. This suggests that SIDE successfully captures a substantial number of moderately similar images, enhancing the diversity and relevance of the extracted data.
  • LIPIPS: SIDE exhibits a lower LIPIPS of 0.809 compared to the baseline’s 0.892. The difference is less pronounced. This is because the primary purpose of LIPIPS is to evaluate perceptual differences between images, rather than structural or training-specific similarity. It measures how humans perceive image similarity based on deep features from pre-trained networks, making it effective for detecting perceptual distortions or quality changes but not for identifying whether an image was directly memorized from the training data.
Visual examples Figure 8 showcases examples of extracted images using the SIDE method alongside the corresponding training images. The visual comparison underscores the high degree of similarity achieved by SIDE, affirming its efficacy in extracting memorized training data.
Figure 8: Examples of Extracted Images using SIDE Method. Top Row: Corresponding training images from the Laion-2B subset. Bottom Row: Images generated by querying the fine-tuned Stable Diffusion model with random strings using the SIDE method. The high visual similarity demonstrates the method’s effectiveness in extracting training data.
The Extended SIDE method effectively extracts training data from fine-tuned DPMs by exploiting the memorization induced by random strings during the fine-tuning process. The quantitative results demonstrate that SIDE significantly enhances the extracted images’ structural and mid-range similarity compared to the baseline. This indicates that SIDE can reliably reproduce specific training images by leveraging the unique associations between random strings and their corresponding images.